The distance between two successive minima of a transverse wave is 2.72 m. Five crests of the wave pass a given point along the direction of travel every 16.1 s.
(a) Find the frequency of the wave.
_________ Hz

(b) Find the wave speed.
_________ m/s
Answer #1

wavelengh \(=2.72 \mathrm{~m}\)

(A) frequency of the wave \(\mathrm{f}=\frac{1}{T}=\frac{1}{16.1 / 5}=0.3105 \mathrm{~Hz}\)

\((B)\)

speed of the wave

\(\mathrm{v}=\mathrm{f} \lambda\)

\(\mathrm{v}=0.3105 \times 2.72=0.8447 \mathrm{~m}\)